# Quantum emitters in silicon: Purcell-enhanced lifetime modulation as a probe of local refractive index changes

### v2

@@ III Advantage of long-lived emitters like T-centers
One of the key technical challenges in high-$Q$ cavity-enhanced sensing is the requirement for very high time resolution in TCSPC systems, especially when using fast-emitting fluorophores like organic dyes ($\tau\sim 3.5$ ns). In those cases, small relative lifetime changes (e.g., 0.1–0.3%) correspond to absolute shifts in the range of 3–10 ps — demanding the use of specialized ultrafast detection equipment. However, T-centers in silicon offer a unique advantage due to their long natural lifetimes, often in the microsecond range. For instance, a T-center with a cavity-unaffected (uncoupled) lifetime of 1 $\mu$[14], when coupled to a photonic cavity with a moderate Purcell factor (e.g., $F_{P}=5$), will exhibit a reduced lifetime of about 200 ns. In this regime, even a 0.1% lifetime change corresponds to a shift of 200 ps — easily detectable with standard TCSPC systems. This greatly relaxes the instrumentation requirements and enables detection of ultra-small refractive index changes, making T-centers a highly attractive emitter platform for integrated biosensing using cavity-enhanced lifetime detection. Figure 3 shows a comparison of effective lifetimes $\tau_{\mathrm{eff}}$ of a typical fluorophore (3.5 ns intrinsic) and a T-center with an uncoupled lifetime of 1 $\mu$s, as a function of Purcell factor. T-centers maintain a large $\tau_{\mathrm{eff}}$ ($\sim$1 ns) even under high Purcell enhancement ($F_{P}=10^{3}$), easing TCSPC requirements.
Figure 3: Comparison of effective lifetimes $\tau_{\mathrm{eff}}$ of a typical fluorophore ($\sim$3.5 ns intrinsic) and a T-center with an uncoupled lifetime of 1 $\mu$s, as a function of Purcell factor.
@@
Table 3: Estimated lifetime and $\Delta\tau$ requirements for fluorescent dyes (3.5 ns) and T-centers (1 $\mu$s) under various $Q$ values assuming off-resonance sensing.
@@
$10^{5}$ Est. $F_{P}$: $\sim$10 $\tau_{\mathrm{eff}}$: $\sim$0.35 ns $\Delta\tau$ (3%): $\sim$10.5 ps Required TCSPC: $\sim$10–20 ps (high-end) Est. $F_{P}$: $\sim$10 $\tau_{\mathrm{eff}}$: $\sim$100 ns $\Delta\tau$ (3%): $\sim$3 ns Required TCSPC: $\sim$2–5 ns (standard)
$10^{6}$ Est. $F_{P}$: $\sim$50 $\tau_{\mathrm{eff}}$: $\sim$0.07 ns $\Delta\tau$ (3%): $\sim$2.1 ps Required TCSPC: $<$10 ps (cutting-edge) Est. $F_{P}$: $\sim$50 $\tau_{\mathrm{eff}}$: $\sim$20 ns $\Delta\tau$ (3%): $\sim$600 ps Required TCSPC: $\sim$500 ps–1 ns (standard)
$10^{7}$ Est. $F_{P}$: $\sim$100 $\tau_{\mathrm{eff}}$: $\sim$0.035 ns $\Delta\tau$ (3%): $\sim$1.0 ps Required TCSPC: $<$5 ps (advanced only) Est. $F_{P}$: $\sim$100 $\tau_{\mathrm{eff}}$: $\sim$10 ns $\Delta\tau$ (3%): $\sim$300 ps Required TCSPC: $\sim$200–500 ps (standard)
As shown in Table 3, fluorescent dyes require very high timing resolution (20 ps or better) for high-$Q$ cavities, making them challenging to use without ultrafast detection systems. In contrast, T-centers, even when strongly coupled to high-$Q$ cavities, maintain effective lifetimes in the tens to hundreds of nanoseconds. This translates into much larger absolute lifetime changes, which can be resolved using standard TCSPC systems with time resolution in the hundreds of picoseconds to a few nanoseconds. This distinction makes T-centers a highly attractive choice for scalable and practical high-sensitivity sensing.
@@ IV Comparison with other approaches based on resonance wavelength shifts
In addition, it is worth comparing the proposed approach with a commercially available alternative. Commercially available platforms such as Delta Life Science’s inQuiQ system [24] utilize microring resonator (MRR) structures to detect analyte-induced refractive index changes via resonance wavelength shifts, achieving state-of-the-art sensitivity with a baseline noise level of $\sim$0.01 RU (corresponding to $\sim 1\times 10^{-8}$ RIU) at a 1 Hz readout. According to publicly available information, this sensitivity is achieved using arrays of moderately high-$Q$ MRRs ($Q\approx 10^{5}$), operated near critical coupling to maximize the resonance shift per refractive index change $\Delta n$, and supported by active temperature stabilization and signal averaging to suppress baseline drift and noise. While highly effective, this approach relies on precise spectral tracking, which imposes integration challenges and demands high-resolution optical readout. It is unlikely that such systems employ on-chip integrated spectrometers, as achieving sub-picometer resolution with current CMOS-compatible technologies remains unfeasible; recent integrated spectrometer designs offer resolutions on the order of $\sim$100 pm [25], which is several orders of magnitude too coarse to resolve the required spectral shifts. In contrast, our platform employs a fundamentally different transduction mechanism based on TCSPC to monitor lifetime shifts of embedded quantum emitters in silicon, directly modulated by changes in the LDOS. Importantly, this lifetime-based approach enables comparable sensitivity with photonic cavities having the same $Q$ values ($10^{5}$$10^{6}$), significantly relaxing fabrication requirements. This makes our approach more scalable and robust, with inherent advantages in CMOS compatibility, noise resilience, and full on-chip integration without the need for spectral scanning or external tunable sources.
@@ V Physical basis of the signal transduction amplifier
The above discussions indicate that this approach may be suitable for use without any effective signal transduction amplifier of the local refractive index change signal (such as an aptamer-functionalized hydrogel (aptagel) layer [26]), as it offers LODs on the order of $10^{-7}$$10^{-9}$ RIU, which is already potentially sufficient for single-molecule detection. However, it is nonetheless interesting to analyze the role of such a mechanical-optical amplifier and to visualize its potential impact on sensing performance. Eq. (13) can be rearranged as follows:
@@
Eq. (19) expresses that, for a given cavity (i.e., fixed $Q$ and $n_{\text{eff}}$), the relationship between refractive index changes $\Delta n$ and the measurable lifetime shift $\Delta\tau/\tau_{0}$ is fixed at the optimal operating point. Therefore, if one increases the actual refractive index shift $\Delta n$ — for example, due to aptagel collapse or shrinkage upon analyte binding — then the same sensing response (i.e., the same slope sensitivity) can be achieved with a higher $\Delta\tau/\tau_{0}$. In other words, while the aptagel does not alter the fundamental sensitivity of the photonic cavity (as governed by $Q$ and $n_{\text{eff}}$), it significantly increases the local refractive index change $\Delta n$ upon analyte binding. This, in turn, enables a proportional increase in the required measurable lifetime change $\Delta\tau/\tau_{0}$, thereby relaxing the timing resolution demands of the TCSPC detection system. In this way, the aptagel acts as an effective mechanical-optical amplifier of the analyte signal. To illustrate this concept, we can consider the following example: Without an aptagel, the binding of an analyte may induce a small refractive index change of approximately $1.2\times 10^{-7}$ RIU. For a photonic cavity with a quality factor $Q=10^{6}$, this corresponds to a relative lifetime change $\Delta\tau/\tau_{0}$ of about 0.3%. Detecting such a small change would require a time resolution of around 60 picoseconds (see Table 3) — a capability available in high-end TCSPC systems. In contrast, when a responsive aptagel is employed, the same analyte binding event could lead to a significantly larger refractive index change — for example, $1.2\times 10^{-6}$ RIU. This results in a relative lifetime change of approximately 3%, which would translate to a TCSPC absolute resolution requirement of $\sim$600 picoseconds. Such changes are much more easily detectable and can be resolved using standard, lower-cost TCSPC instrumentation. In this way, the aptagel does not increase the intrinsic sensitivity of the photonic cavity but instead relaxes the detection requirements by amplifying the local refractive index change resulting from analyte binding.
@@
In addition, based on some realistic assumptions, we have attempted to “visualize” the effect of the aptagel in comparison to direct molecular binding to the sensor surface. When a single molecule binds directly to the surface of a SiO2 cavity, it perturbs the cavity’s optical mode through its local refractive index contribution. This perturbation is relatively small, as the molecule interacts only with the evanescent tail of the cavity mode. Let us assume the following: the refractive index of the molecule is $n_{\text{mol}}\approx 1.45$, the background medium has a refractive index of $n_{\text{bg}}\approx 1.33$, the volume of a single molecule is $V_{\text{mol}}\approx 10^{2}$ nm3, and the mode volume of the cavity is $V_{\text{mode}}\approx 10^{5}$ nm3. Then, the effective refractive index perturbation can be estimated as $\Delta n^{\text{direct}}=(n_{\text{mol}}-n_{\text{bg}})\cdot\left(V_{\text{mol}}/V_{\text{mode}}\right)\cdot f_{\text{field}}(\vec{r})$, where $f_{\text{field}}(\vec{r})$ is the normalized field intensity at the molecule’s position $\vec{r}$, relative to the mode field maximum. Assuming $f_{\text{field}}(\vec{r})=10^{-3}$, the resulting $\Delta n^{\text{direct}}\approx 1.2\times 10^{-7}$ RIU, which is in good agreement with our earlier estimation based on Ref. [23].
@@ VI Conclusion
Additionally, we show that incorporating functional layers such as aptagels can act as signal transduction amplifiers, further increasing the effective refractive index shift and thereby reducing the demands on TCSPC detection performance.
In summary, this work establishes a fundamentally different and technologically viable sensing modality — one that combines the robustness of quantum emitters with the scalability of silicon photonics and the practicality of lifetime-based detection. Although room-temperature operation of quantum emitters in silicon has yet to be demonstrated, the proposed platform offers a clear pathway toward next-generation, on-chip sensors with quantum-level sensitivity, operable at room temperature and without reliance on spectral resolution.
@@ Acknowledgements.
The data that support the findings of this article are included within the article.

### v4

@@ III Advantage of long-lived emitters like T-centers
[…]
@@
[…]
@@
$10^{5}$
Est. $F_{P}$: $\sim$10
$\tau_{\mathrm{eff}}$: $\sim$0.35 ns
$\Delta\tau$ (3%): $\sim$10.5 ps
Required TCSPC: $\sim$10–20 ps (high-end)
Est. $F_{P}$: $\sim$10
$\tau_{\mathrm{eff}}$: $\sim$100 ns
$\Delta\tau$ (3%): $\sim$3 ns
Required TCSPC: $\sim$2–5 ns (standard)
$10^{6}$
Est. $F_{P}$: $\sim$50
$\tau_{\mathrm{eff}}$: $\sim$0.07 ns
$\Delta\tau$ (3%): $\sim$2.1 ps
Required TCSPC: <<10 ps (cutting-edge)
Est. $F_{P}$: $\sim$50
$\tau_{\mathrm{eff}}$: $\sim$20 ns
$\Delta\tau$ (3%): $\sim$600 ps
Required TCSPC: $\sim$500 ps–1 ns (standard)
$10^{7}$
Est. $F_{P}$: $\sim$100
$\tau_{\mathrm{eff}}$: $\sim$0.035 ns
$\Delta\tau$ (3%): $\sim$1.0 ps
Required TCSPC: <<5 ps (advanced only)
Est. $F_{P}$: $\sim$100
$\tau_{\mathrm{eff}}$: $\sim$10 ns
$\Delta\tau$ (3%): $\sim$300 ps
Required TCSPC: $\sim$200–500 ps (standard)
As shown in Table 3, fluorescent dyes require very high timing resolution (20 ps or better) for high-$Q$ cavities, making them challenging to use without ultrafast detection systems. In contrast, T-centers, even when strongly coupled to high-$Q$ cavities, maintain effective lifetimes in the tens to hundreds of nanoseconds. This translates into much larger absolute lifetime changes, which can be resolved using standard TCSPC systems with time resolution in the hundreds of picoseconds to a few nanoseconds. This distinction makes T-centers a highly attractive choice for scalable and practical high-sensitivity sensing. It should be noted that the “Required TCSPC” values in Table III are provided as illustrative benchmarks of commercially available temporal resolution and should not be interpreted as strict physical limits. In practice, the detectability of small fractional lifetime changes is determined primarily by photon statistics and signal-to-noise ratio. Coarser binning may be employed to improve photon counts per bin without significantly degrading the ability to resolve changes in decay slope.
@@ IV Comparison with other approaches based on resonance wavelength shifts
In addition, it is worth comparing the proposed approach with a commercially available alternative. Commercially available platforms such as Delta Life Science’s inQuiQ system [28] utilize microring resonator (MRR) structures to detect analyte-induced refractive index changes via resonance wavelength shifts, achieving state-of-the-art sensitivity with a baseline noise level of $\sim$0.01 RU (corresponding to $\sim 1\times 10^{-8}$ RIU) at a 1 Hz readout (for ease of comparison, in some cases the calculated refractive index change (in RIU) is converted to resonance units (RU) — a standard metric in SPR and other sensing techniques — using the relation $1\,\mathrm{RU}=10^{-6}\,\mathrm{RIU}$). According to publicly available information, this sensitivity is achieved using arrays of moderately high-$Q$ MRRs ($Q\approx 10^{5}$), operated near critical coupling to maximize the resonance shift per refractive index change $\Delta n$, and supported by active temperature stabilization and signal averaging to suppress baseline drift and noise. While highly effective, this approach relies on precise spectral tracking, which imposes integration challenges and demands high-resolution optical readout. It is unlikely that such systems employ on-chip integrated spectrometers, as achieving sub-picometer resolution with current CMOS-compatible technologies remains unfeasible; recent integrated spectrometer designs offer resolutions on the order of $\sim$100 pm [29], which is several orders of magnitude too coarse to resolve the required spectral shifts. In contrast, our platform employs a fundamentally different transduction mechanism based on TCSPC to monitor lifetime shifts of embedded quantum emitters in silicon, directly modulated by changes in the LDOS. Importantly, this lifetime-based approach enables comparable sensitivity with photonic cavities having the same $Q$ values ($10^{5}$$10^{6}$), significantly relaxing fabrication requirements. This makes our approach more scalable and robust, with inherent advantages in CMOS compatibility, noise resilience, and full on-chip integration without the need for spectral scanning or external tunable sources.
@@ V Physical basis of the signal transduction amplifier
The above discussions indicate that this approach may be suitable for use without any effective signal transduction amplifier of the local refractive index change signal (such as an aptamer-functionalized hydrogel (aptagel) layer [30], by which the interaction areas schematically depicted in Figure 1 can be covered), as it offers LODs on the order of $10^{-7}$$10^{-9}$ RIU, which is already potentially sufficient for single-molecule detection. However, it is nonetheless interesting to analyze the role of such a mechanical-optical amplifier and to visualize its potential impact on sensing performance.
For clarity, we note here that an aptagel is a hydrogel functionalized with specific aptamers (short single-stranded DNA or RNA sequences) that selectively bind target molecules. Upon binding, the aptamer undergoes a conformational change, which can trigger a macroscopic collapse or shrinkage of the hydrogel network (aptagel collapse). This collapse increases the local polymer density (polymer volume fraction), thereby increasing the local refractive index. In this way, the aptagel acts as a signal amplifier: a single molecular binding event can produce a much larger effective $\Delta n$ than direct binding of the molecule to the cavity surface. This differs from multiple identical molecules binding simultaneously, because here the amplification is not merely additive but comes from a cooperative volumetric change in the material.
Eq. (13) can be rearranged as follows:
@@
Eq. (21) expresses that, for a given cavity (i.e., fixed $Q$ and $n_{\text{eff}}$), the relationship between refractive index changes $\Delta n$ and the measurable lifetime shift $\Delta\tau/\tau_{0}$ is fixed at the optimal operating point. Therefore, if one increases the actual refractive index shift $\Delta n$ — for example, due to aptagel collapse upon analyte binding — then the same sensing response (i.e., the same slope sensitivity) can be achieved with a higher $\Delta\tau/\tau_{0}$. In other words, while the aptagel does not alter the fundamental sensitivity of the photonic cavity (as governed by $Q$ and $n_{\text{eff}}$), it significantly increases the local refractive index change $\Delta n$ upon analyte binding. This, in turn, enables a proportional increase in the required measurable lifetime change $\Delta\tau/\tau_{0}$, thereby relaxing the timing resolution demands of the TCSPC detection system.
To illustrate this concept, consider the following example: Without an aptagel, the binding of an analyte may induce a small refractive index change of approximately $1.2\times 10^{-7}$ RIU. For a photonic cavity with a quality factor $Q=10^{6}$, this corresponds to a relative lifetime change $\Delta\tau/\tau_{0}$ of about 0.3%. Detecting such a small change would require a time resolution of around 60 ps (see Table 3) — a capability available in high-end TCSPC systems. In contrast, when a responsive aptagel is employed, the same analyte binding event could lead to a significantly larger refractive index change — for example, $1.2\times 10^{-6}$ RIU. This results in a relative lifetime change of approximately 3%, which would translate to a TCSPC absolute resolution requirement of $\sim$600 ps. Such changes are much more easily detectable and can be resolved using standard, lower-cost TCSPC instrumentation. In this way, the aptagel does not increase the intrinsic sensitivity of the photonic cavity but instead relaxes the detection requirements by amplifying the local refractive index change resulting from analyte binding.
@@
In addition, based on some realistic assumptions, we have attempted to visualize the effect of the aptagel in comparison to direct molecular binding to the sensor surface. When a single molecule binds directly to the surface of a SiO2 cavity, it perturbs the cavity’s optical mode through its local refractive index contribution. This perturbation is relatively small, as the molecule interacts only with the evanescent tail of the cavity mode. Let us assume the following: the refractive index of the molecule is $n_{\text{mol}}\approx 1.45$, the background medium has a refractive index of $n_{\text{bg}}\approx 1.33$, the volume of a single molecule is $V_{\text{mol}}\approx 10^{2}$ nm3, and the mode volume of the cavity is $V_{\text{mode}}\approx 10^{5}$ nm3. Then, the effective refractive index perturbation can be estimated as $\Delta n^{\text{direct}}=(n_{\text{mol}}-n_{\text{bg}})\cdot\left(V_{\text{mol}}/V_{\text{mode}}\right)\cdot f_{\text{field}}(\vec{r})$, where $f_{\text{field}}(\vec{r})$ is the normalized field intensity at the molecule’s position $\vec{r}$, relative to the mode field maximum. Assuming $f_{\text{field}}(\vec{r})=10^{-3}$, the resulting $\Delta n^{\text{direct}}\approx 1.2\times 10^{-7}$ RIU, which is in good agreement with our earlier estimation based on Ref. [27].
@@ VI Discussion and Conclusion
Figure 7: Possible application pathway of the proposed sensing platform for single-molecule interaction studies in personalized medicine.
In terms of practical implementation, the sensing concept is directly compatible with silicon chips integrating single-photon avalanche diodes (SPADs), enabling label-free, on-chip lifetime measurements with high timing precision. Quality factor stability and cavity alignment — often critical challenges in high-$Q$ resonance-shift sensors — are less stringent in our approach due to the off-resonance operating regime, which maintains sensitivity even in the presence of moderate cavity detuning. According to Eq. (12) and as shown in Table 1, the sensitivity of the proposed lifetime-based sensing scheme scales linearly with the cavity quality factor $Q$. Consequently, any relative variation in $Q$ directly translates into an equivalent relative change in sensitivity; for example, a $\pm 5\%$ fluctuation in $Q$ will result in approximately a $\pm 5\%$ change in sensitivity. This relationship underscores the importance of maintaining $Q$-factor stability in practical implementations, particularly when targeting ultra–low detection limits. Moreover, multiplexed layouts (Fig. 1) can allow parallel operation across multiple sensing channels, increasing throughput and enabling multi-analyte detection.
Figure 7 outlines a possible application pathway in which the proposed platform serves as a basis for single-molecule interaction studies in personalized medicine. Here, a CMOS photonic integrated circuit (PIC) with integrated SPADs and interaction areas performs lifetime-based sensing. Real-time decay dynamics are processed via AI-assisted algorithms to extract binding events, and the resulting information can directly feed into clinical decision support systems, for example, to optimize immunosuppressive drug dosing in transplant patients.
While full TCSPC decay acquisition can take longer than the timescales of some single-molecule binding/unbinding events, detection speed can be improved by using high-brightness emitters, optimized photon extraction, and fast analysis techniques (e.g., phasor plots, running-window fits, or machine-learning-based event detection) that operate on partial decay data. Such approaches enable practical lifetime-based sensing even in dynamic environments and, while may not achieve instantaneous detection like resonance shift sensors, it can still operate effectively within the binding/unbinding timescales relevant to many biosensing applications. Furthermore, employing a pulsed laser with a MHz repetition rate together with high-speed acquisition and post-processing hardware ensures that TCSPC does not present a limitation for the proposed detection scheme.
Additionally, incorporating functional layers such as aptagels can act as signal transduction amplifiers, further increasing the effective refractive index shift and thereby reducing the demands on TCSPC detection performance.
In summary, this work establishes a fundamentally different and technologically viable sensing modality — one that combines the robustness of quantum emitters with the practicality of lifetime-based detection, compatibility with SPAD arrays for label-free operation, and seamless integration into scalable PIC platforms. While silicon has been used here as a demonstrator owing to its CMOS compatibility and mature fabrication ecosystem, the generic nature of the proposed approach makes it readily adaptable to other PIC materials in which room-temperature operation of quantum emitters has already been experimentally demonstrated, such as diamond, silicon carbide, and silicon nitride. This versatility opens a clear pathway toward next-generation, on-chip sensors with quantum-enabled sensitivity, operable at room temperature and without reliance on spectral resolution.
Author Contributions
Y.M. conceived the idea, performed the calculations, developed the Python scripts, and wrote the original draft of the manuscript. A.L. proposed the model of the averaged fluorescence decay (Eq. (18)), assisted with the calculations, and contributed to manuscript writing. All authors have read and approved the final version of the manuscript.
@@ Acknowledgements.
This study did not generate new experimental data. All Python scripts used for analytical calculations and figure generation, as well as a supporting PDF document detailing LDR calculation, are available at the OSF repository (under the Files tab): https://doi.org/10.17605/OSF.IO/QDYNX.